# DP-GTR: Differentially Private Prompt Protection via Group Text Rewriting

### v1

@@ Abstract
Prompt privacy is crucial, especially when using online large language models (LLMs), due to the sensitive information often contained within prompts. While LLMs can enhance prompt privacy through text rewriting, existing methods primarily focus on document-level rewriting, neglecting the rich, multi-granular representations of text. This limitation restricts LLM utilization to specific tasks, overlooking their generalization and in-context learning capabilities, thus hindering practical application. To address this gap, we introduce DP-GTR, a novel three-stage framework that leverages local differential privacy (DP) and the composition theorem via group text rewriting. DP-GTR is the first framework to integrate both document-level and word-level information while exploiting in-context learning to simultaneously improve privacy and utility, effectively bridging local and global DP mechanisms at the individual data point level. Experiments on CommonSense QA and DocVQA demonstrate that DP-GTR outperforms existing approaches, achieving a superior privacy-utility trade-off. Furthermore, our framework is compatible with existing rewriting techniques, serving as a plug-in to enhance privacy protection. Our code is publicly available at github.com for reproducibility.
@@
Mingchen Lithanks: Corresponding author: mingchen.li@unt.edu, Heng Fan, Song Fu, Junhua Ding, and Yunhe Feng Department of Computer Science and Engineering University of North Texas {mingchen.li, heng.fan, song.fu, junhua.ding, yunhe.feng}@unt.edu
@@ 1 Introduction
The rise of LLMs in natural language processing has spurred research into differential privacy (DP) techniques to mitigate the risk of sensitive information leakage (Abadi et al., 2016; Wu et al., 2023; Tang et al., 2023). While DP, the gold standard for computational privacy, has seen broad adoption in machine learning, existing text-based DP methods face significant challenges. These methods generally fall into four categories: training-based optimizations (e.g., DP-SGD (Abadi et al., 2016)), embedding perturbations (Feyisetan et al., 2020), document-level paraphrasing (Mattern et al., 2022), and in-context learning (ICL) enhancements (Wu et al., 2023). However, training-based approaches are computationally expensive, embedding perturbations can compromise semantic coherence, and ICL often neglects client-side prompt privacy.
Document paraphrasing offers a promising balance between privacy and utility. State-of-the-art methods achieve differentially private next-token generation using the exponential mechanism (EM) (McSherry and Talwar, 2007; Carvalho et al., 2023), replacing the standard softmax. Initial work employed decoder-only models like fine-tuned GPT-2 (Mattern et al., 2022), progressing to encoder-decoder (Igamberdiev and Habernal, 2023) and encoder-only architectures (e.g., BART, RoBERTa) (Meisenbacher et al., 2024b). DP-Prompt (Utpala et al., 2023) leverages prompt learning for zero-shot paraphrasing, and recent advancements combine DP post-processing with adversarial fine-tuning (Meisenbacher and Matthes, 2024). A critical limitation, however, persists: the lack of fine-grained control over the privacy-utility trade-off.
@@ 2 Related Work
Global vs. Local DP: Differential privacy text sanitization methods are classified into Global Differential Privacy (Global-DP) and Differential Privacy (Local-DP) based on where the privacy mechanism is applied. In Global-DP, data is aggregated centrally before applying the privacy mechanism, while methods like DP-SGD use differentially private optimization techniques for training text models.  (Abadi et al., 2016; Ponomareva et al., 2022; Feyisetan et al., 2020). DP-ICL operates within a “sample-and-aggregate” framework by perturbing the embedding and vocabulary selected for release  (Wu et al., 2023). In contrast, Local-DP incorporates the differential privacy mechanism before data reaches the centralized processor, typically affording stronger privacy protection  (Duchi et al., 2013; Feyisetan et al., 2020).
@@
Local-DP: Private document release methods are categorized into three tiers based on where noise is added: word-level, sentence-level, and document-level. At the word level, noise is added to word embeddings, and the perturbed vectors are then mapped to the nearest vocabulary word (Feyisetan et al., 2020; Xu et al., 2020; Yue et al., 2021).  Carvalho et al. (2023) employ the exponential mechanism for token selection, while  Chen et al. (2023a) propose customized token mappings for individual words. Moreover,  Meisenbacher et al.’s (2024a) study generates multiple candidate perturbations using various word embedding models, and  Mattern et al. (2022) highlight that word-level approaches inherently lack contextual information. Similarly, sentence-level methods inject noise into sentence embeddings  (Reimers, 2019; Meehan et al., 2022).
@@
Document-level: At the document level, paraphrasing technologies are grouped into three categories based on model architecture. Mattern et al. (2022) uses a decoder-only fine-tuned GPT-2 model. Later work adopts encoder-decoder models, such as a BART-based approach (Lewis, 2019) with sensitivity clipping via thresholding and pruning (Igamberdiev and Habernal, 2023). Encoder-only methods, like DP-MLM (Meisenbacher et al., 2024b), use a RoBERTa-based masked language model for fine-tuning. These methods require fine-tuning. In contrast, DP-Prompt Utpala et al.’s (2023) introduces a zero-shot prompt learning paradigm using black-box LLMs, and Meisenbacher and Matthes (2024) employ post-processing and adversarial fine-tuning to enhance rewriting.
@@
DP in ICL: The primary concern with applying DP in ICL is that LLMs are not inherently secure, potentially exposing sensitive context. Wu et al. (2023)’s DP-ICL perturbs embeddings and extracts keywords from subsampled datasets, while Tang et al. (2023) incorporate label information. Zheng et al. (2024) employ $k$-RR (Wang et al., 2017) to generate ICL answers, and Gao et al. (2024) aggregate next-token predictions from dataset shards. All the approaches follow a “sample-and-aggregate” framework (Nissim et al., 2007a), partitioning the data and applying private aggregation.
@@ 3 Preliminaries
Prior work  (Utpala et al., 2023; Mattern et al., 2022) has shown the equivalence between this softmax selection process and the Exponential Mechanism (EM) of differential privacy, where the utility function corresponds to the logits. Assuming $LM$ is not pre-trained on the distribution of the data being protected, and that logits $u_{v}$ are clipped to $[b_{\min},b_{\max}]$, generating $n$ tokens at temperature $T$ provides a $(\frac{2n(b_{\max}-b_{\min})}{T})$-local DP (LDP) guarantee. This derives from the fact that each token selection, with a maximum logit difference of $(b_{\max}-b_{\min})$, incurs a privacy loss of $\frac{2(b_{\max}-b_{\min})}{T}$. Sequential composition over $n$ tokens then yields the stated LDP bound for a single document paraphrase. Logit clipping and EM sampling ensure the generated sequence respects a well-defined pure LDP budget. See Appendix A for Algorithm 2 and proof.
@@ 4.2 Stage-1: Group Text Rewriting
DP-GTR employs group text rewriting to achieve finer-grained control over privacy and utility compared to document-level methods. Effective prompt sanitization requires considering both document-level context for overall meaning and word-level information for protecting sensitive terms. While LLMs excel with contextual input, directly using the original prompt for in-context learning compromises privacy.
Group text rewriting addresses this by generating a local paraphrased text database, effectively mitigating the limitations of both document-level rewriting and the absence of suitable contextual information. This database, consisting of multiple rewritten versions of the prompt, serves several key purposes. First, it provides richer information than a single rewrite, capturing diverse facets of the original prompt. Second, aggregating word counts across the rewrites facilitates the identification of shared, potentially sensitive keywords. More importantly, the group enables more aggressive post-processing and additional DP mechanisms, which, while not reducing the formal $\epsilon$-DP bound, further mitigate the risk of sensitive information disclosure by eliminating sensitive identifiers and limiting real-world exposure. Specifically, generating a paraphrased group $\mathcal{P}$ of $m$ documents, each with $n$ tokens, incurs an $mn\epsilon$-DP privacy budget.
@@ 4.3.2 Consensus-aware privacy protection
13:     return  JointEM-TopK          (private_keywords,$\epsilon$, $K$)
@@ 4.4 Stage-3: Privacy-Preserving Prompting
Privacy-Preserving Prompt Template Refer to the following question to generate a new question: <$P_{low}$> Avoid using the following tokens: <$w_{1},w_{2},...,w_{k}$>
@@ 5.2 Experiment Setups
DP-GTR Settings. We use GPT-3.5 (black-box) and Llama-3.1-8B (white-box) as underlying models for DP-GTR. Both the number of group text rewritings and private keywords are set to 10. For private keyword release in Stage-2, we implement a non-DP post-processing method (DP-GTR-NDP) and a differentially private JointEM mechanism (DP-GTR-JEM). In Stage-1, paraphrasing is controlled by temperature (black-box) and privacy budget $\epsilon$ (white-box), with nine values tested: $T\in\{0.1,0.15,\dots,1.5\}$. Corresponding $\epsilon$ values for the white-box model are calculated based on temperature and pre-clipped sensitivity (see Appendix Section A.2). Stage-3 prompt generation uses a temperature of 0 (or equivalently low) without a DP mechanism.
@@
(a) Rouge1 (GPT-3.5)
(b) Rouge1 (Llama)
(c) RougeL (Llama)
(d) RougeL (Llama)
(e) BLEU (GPT-3.5)
(f) BLEU (Llama)
@@ 5.3 Results on Open-answer VQA Dataset
Methods Question (Privacy Leakage) Answer (Utility: Rouge1)
Rouge1 RougeL BLEU Value
@@
Methods Question (Privacy Leakage) Answer (Utility: Accuracy)
Rouge1 RougeL BLEU Value
@@
5.4 Results on Close-answer CSQA Dataset
5.5 Non-Uniform Rewriting Strategies
@@ 6 Conclusion
This paper proposes DP-GTR, a novel three-stage local differential privacy (LDP) framework that leverages differentially private paraphrasing and the composition theorem through group text rewriting to enhance the privacy-utility trade-off. DP-GTR is the first approach, to our knowledge, to apply in-context learning for LDP prompt privatization and to connect global and local DP mechanisms via grouped paraphrased text. Furthermore, our framework is generalizable and compatible with any existing paraphrasing technique. Evaluations on open- and closed-answer QA datasets (DocVQA and Commonsense QA), simulating real-world LLM application scenarios, demonstrate that DP-GTR achieves a significantly superior privacy-utility trade-off compared to existing state-of-the-art methods. With the rapidly increasing adoption of LLMs, DP-GTR provides a practical, robust, and readily deployable solution for mitigating the risk of user prompt privacy leakage.
@@ 7 Limitations
Appendix A Appendix
A.1 DP-guaranteed Paraphrasing Proof
Proof. Let $D$ and $D^{\prime}$ be any two documents, and $\mathbf{u}$ and $\mathbf{u}^{\prime}\in\mathbb{R}^{|V|}$ be their corresponding logits and $\,[b_{min},\,b_{max}]$ is the minium and maximum value of the logit. Let $v\in V$, and $i$ be its index, with $u_{i}$ being its corresponding logit. We then have that (Utpala et al., 2023),
$\frac{\text{Pr}[M(D)=v]}{\text{Pr}[M(D^{\prime})=v]}&=\frac{\frac{\exp(\frac{u_{i}}{T})}{\sum_{j=1}^{|V|}\exp(\frac{u_{j}}{T})}}{\frac{\exp(\frac{u_{i}^{\prime}}{T})}{\sum_{j=1}^{|V|}\exp(\frac{u_{j}^{\prime}}{T})}}\\ &=\frac{\exp(\frac{u_{i}}{T})}{\exp(\frac{u_{i}^{\prime}}{T})}\frac{\sum_{j=1}^{|V|}\exp(\frac{u_{j}^{\prime}}{T})}{\sum_{j=1}^{|V|}\exp(\frac{u_{j}}{T})}\\ &=\exp\left(\frac{u_{i}-u_{i}^{\prime}}{T}\right)\frac{\sum_{j=1}^{|V|}\exp(\frac{u_{j}^{\prime}}{T})}{\sum_{j=1}^{|V|}\exp(\frac{u_{j}}{T})}\\ &\leq\exp\left(\frac{b_{max}-b_{min}}{T}\right)\exp\left(\frac{b_{max}-b_{min}}{T}\right)\\ &\leq\exp\left(\frac{2(b_{max}-b_{min})}{T}\right).\end{aligned}$

### v2

@@ Abstract
Prompt privacy is crucial, especially when using online large language models (LLMs), due to the sensitive information often contained within prompts. While LLMs can enhance prompt privacy through text rewriting, existing methods primarily focus on document-level rewriting, neglecting the rich, multi-granular representations of text. This limitation restricts LLM utilization to specific tasks, overlooking their generalization and in-context learning capabilities, thus hindering practical application. To address this gap, we introduce DP-GTR, a novel three-stage framework that leverages local differential privacy (DP) and the composition theorem via group text rewriting. DP-GTR is the first framework to integrate both document-level and word-level information while exploiting in-context learning to simultaneously improve privacy and utility, effectively bridging local and global DP mechanisms at the individual data point level. Experiments on CommonSense QA and DocVQA demonstrate that DP-GTR outperforms existing approaches, achieving a superior privacy-utility trade-off. Furthermore, our framework is compatible with existing rewriting techniques, serving as a plug-in to enhance privacy protection. Our code is publicly available at github.com/ResponsibleAILab/DP-GTR.
@@
[…]
@@ 1 Introduction
The rise of LLMs in natural language processing has catalyzed an urgent research focus on their security and privacy Yao et al. (2024); Peng et al. (2025); Luo et al. (2025), including investigations into differential privacy (DP) techniques to mitigate the risk of sensitive information leakage (Edemacu and Wu, 2025; Abadi et al., 2016; Wu et al., 2023; Tang et al., 2023). While DP, the gold standard for computational privacy, has seen broad adoption in machine learning, existing text-based DP methods face significant challenges. These methods generally fall into four categories: training-based optimizations (e.g., DP-SGD (Abadi et al., 2016)), embedding perturbations (Feyisetan et al., 2020), document-level paraphrasing (Mattern et al., 2022), and in-context learning (ICL) enhancements (Wu et al., 2023). However, training-based approaches are computationally expensive, embedding perturbations can compromise semantic coherence, and ICL often neglects client-side prompt privacy.
Document paraphrasing offers a promising balance between privacy and utility. State-of-the-art methods achieve differentially private next-token generation using the exponential mechanism (EM) (McSherry and Talwar, 2007; Carvalho et al., 2023), replacing the standard softmax. Initial work employed decoder-only models like fine-tuned GPT-2 (Mattern et al., 2022), progressing to encoder-decoder (Igamberdiev and Habernal, 2023) and encoder-only architectures (e.g., BART, RoBERTa) (Meisenbacher et al., 2024b). DP-Prompt (Utpala et al., 2023) leverages prompt learning for zero-shot paraphrasing, and recent advancements combine DP post-processing with adversarial fine-tuning (Meisenbacher and Matthes, 2024). A critical limitation, however, persists: lack of fine-grained control over the privacy-utility trade-off.
@@ 2 Related Work
Global vs. Local DP: Differential privacy text sanitization methods are classified into Global Differential Privacy (Global-DP) and Local Differential Privacy (Local-DP) based on where the privacy mechanism is applied. In Global-DP, data is aggregated centrally before applying the privacy mechanism, while methods like DP-SGD use differentially private optimization techniques for training text models (Abadi et al., 2016; Ponomareva et al., 2022; Feyisetan et al., 2020). DP-ICL operates within a “sample-and-aggregate” framework by perturbing the embedding and vocabulary selected for release (Wu et al., 2023). In contrast, Local-DP incorporates the differential privacy mechanism before data reaches the centralized processor, typically affording stronger privacy protection (Duchi et al., 2013; Feyisetan et al., 2020).
@@
Local-DP: Private document release methods are categorized into three tiers based on where noise is added: word-level, sentence-level, and document-level. At the word level, noise is added to word embeddings, and the perturbed vectors are then mapped to the nearest vocabulary word (Feyisetan et al., 2020; Xu et al., 2020; Yue et al., 2021).  Carvalho et al. (2023) employ the exponential mechanism for token selection, while  Chen et al. (2023a) propose customized token mappings for individual words. Moreover,  Meisenbacher et al.’s (2024a) study generates multiple candidate perturbations using various word embedding models, and  Mattern et al. (2022) highlights that word-level approaches inherently lack contextual information. Similarly, sentence-level methods inject noise into sentence embeddings  (Reimers, 2019; Meehan et al., 2022).
@@
Document-level: At the document level, paraphrasing technologies are grouped into three categories based on model architecture. Mattern et al. (2022) uses a decoder-only fine-tuned GPT-2 model. Later work adopts encoder-decoder models, such as a BART-based approach (Lewis, 2019) with sensitivity clipping via thresholding and pruning (Igamberdiev and Habernal, 2023). Encoder-only methods, like DP-MLM (Meisenbacher et al., 2024b), use a RoBERTa-based masked language model for fine-tuning. These methods require fine-tuning. In contrast, DP-Prompt (Utpala et al., 2023) introduces a zero-shot prompt learning paradigm using black-box LLMs, and Meisenbacher and Matthes (2024) employs post-processing and adversarial fine-tuning to enhance rewriting.
@@
DP in ICL: The primary concern with applying DP in ICL is that LLMs are not inherently secure, potentially exposing sensitive context. Wu et al. (2023)’s DP-ICL perturbs embeddings and extracts keywords from subsampled datasets, while Tang et al. (2023) incorporates label information. Zheng et al. (2024) employ $k$-$R$ (Wang et al., 2017) to generate ICL answers, and Gao et al. (2024) aggregate next-token predictions from dataset shards. All the approaches follow a “sample-and-aggregate” framework (Nissim et al., 2007a), partitioning the data and applying private aggregation.
@@ 3 Preliminaries
Threat Model. We assume an adversary attacker on the side of the cloud-based model vendor, whose objective is to extract private information (such as personal identifiable information) from the confidential transmission content. The adversary’s access is limited to a customized prompt supplied by the client, but they are free to facilitate their attack.
@@
Prior work  (Utpala et al., 2023; Mattern et al., 2022) has shown the equivalence between this softmax selection process and the Exponential Mechanism (EM) of differential privacy, where the utility function corresponds to the logits. Assuming $LM$ is not pre-trained on the distribution of the data being protected, and that logits $u_{v}$ are clipped to $b_{\min},b_{\max}]$, generating $n$ tokens at temperature $T$ provides a $(\frac{2n(b_{\max}-b_{\min})}{T})$-local DP (LDP) guarantee. This derives from the fact that each token selection, with a maximum logit difference of $(b_{\max}-b_{\min})$, incurs a privacy loss of $\frac{2(b_{\max}-b_{\min})}{T}$. Sequential composition over $n$ tokens then yields the stated LDP bound for a single document paraphrase. Logit clipping and EM sampling ensure the generated sequence respects a well-defined pure LDP budget. See Appendix A for details.
@@ 4.2 Stage-1: Group Text Rewriting
DP-GTR employs group text rewriting to achieve finer-grained control over privacy and utility compared to document-level methods. Effective prompt sanitization requires considering both document-level context for overall meaning and word-level information for protecting sensitive terms. While LLMs excel with contextual input, directly using the original prompt compromises privacy.
Group text rewriting addresses this by generating a local paraphrased text database, effectively mitigating the limitations of both document-level rewriting and the absence of suitable contextual information. This database, consisting of multiple rewritten versions of the prompt, serves several key purposes. First, it provides richer information than a single rewrite, capturing diverse facets of the original prompt. Second, aggregating word counts across the rewrites facilitates the identification of shared, potentially sensitive keywords. More importantly, the group enables more aggressive post-processing and additional DP mechanisms, which, while not reducing the formal $\epsilon$-DP bound, further mitigate the risk of sensitive information disclosure by eliminating sensitive identifiers and limiting real-world exposure. Specifically, generating a paraphrased group $\mathcal{P}$ of $m$ documents, each with $n$ tokens, incurs an $(mn)\epsilon_{1}$-DP privacy budget.
@@ 4.3.2 Consensus-aware privacy protection
[…]
@@ 4.4 Stage-3: Privacy-Preserving Prompting
Privacy-Preserving Prompt Template Refer to the following question to generate a new question: $P_{low}$> Avoid using the following tokens: $w_{1},w_{2},...,w_{k}$>
@@ 5.2 Experiment Setups
(a) Rouge1 (GPT-3.5)
(b) Rouge1 (Llama)
(c) RougeL (GPT-3.5)
(d) RougeL (Llama)
(e) BLEU (GPT-3.5)
(f) BLEU (Llama)
DP-GTR Settings. We use GPT-3.5 (black-box) and Llama-3.1-8B (white-box) as underlying models for DP-GTR. Both the number of group text rewritings and private keywords are set to 10. For private keyword release in Stage-2, we implement a non-DP post-processing method (DP-GTR-NDP) and a differentially private JointEM mechanism (DP-GTR-JEM). In Stage-1, paraphrasing is controlled by temperature (black-box) and privacy budget $\epsilon$ (white-box), with nine values tested: $T\in\{0.1,0.15,\dots,1.5\}$. Corresponding $\epsilon$ values for the white-box model are calculated based on temperature and pre-clipped sensitivity (see Appendix Section B). Stage-3 prompt generation uses a temperature of 0 (or equivalently low) without a DP mechanism.
@@ 5.3 Results on Open-answer VQA Dataset
5.4 Results on Close-answer CSQA Dataset
5.5 Non-Uniform Rewriting Strategies
@@
Methods Question (Privacy Leakage) Answer (Utility)
Rouge1 RougeL BLEU Rouge1
@@
Methods Question (Privacy Leakage) Answer (Utility)
Rouge1 RougeL BLEU Accuracy
@@ 5.6 Generalizable Plug-in Framework
6 Ablation Study
The ablation experiments are conducted on two additional configurations: one with references but without keywords, and another with keywords but without references. The configuration is GPT-3.5 and DocVQA. The results are shown in Tab. 2.
Table 2: Question (Privacy Leakage) and Answer (Utility) ablation study. Subscripts indicate the difference between without Reference/Keywords and Ours-NDP. The privacy and utility metric is Rouge-1 (%) between final prompt and original prompt, and Rouge-1 (%) between query answer and ground-truth answer.
Temperature $T=0.1$ $T=0.25$ $T=0.5$ $T=1.0$
Question (Privacy Leakage: Rouge-1)
Ours-NDP 53.21 53.05 48.91 42.46
w/o Reference 7.74↓45.47 8.12↓44.93 8.45↓40.46 9.44↓33.02
w/o Keywords 50.11↓3.10 49.36↓3.69 48.05↓0.86 47.05↑4.59
Answer (Utility: Rouge-1)
Ours-NDP 55.31 54.51 47.89 34.03
w/o Reference 1.01↓54.30 1.04↓53.47 1.13↓46.76 1.57↓32.46
w/o Keywords 21.61↓33.70 23.08↓31.43 22.76↓25.13 24.58↓9.45
Table 3: The results of the adversarial attack. The Rouge-1 (%) scores measuring question similarity between the original and DP-GTR-processed prompts (Ours), as well as under Static and Adaptive attacks, across different temperatures for both NDP and JEM. Subscripts indicate change relative to ours.
Temperature $T=0.1$ $T=0.25$ $T=0.5$ $T=1.0$
Ours-NDP 53.19 53.04 48.95 42.43
Static Attack 52.27↓0.92 52.37↓0.67 48.31↓0.64 40.21↓2.22
Adaptive Attack 49.87↓3.32 49.54↓3.50 45.01↓3.94 36.93↓5.50
Ours-JEM 53.50 53.77 50.92 46.17
Static Attack 51.87↓1.63 52.67↓1.10 50.50↓0.42 43.68↓2.49
Adaptive Attack 49.60↓3.90 50.48↓3.29 46.90↓4.02 39.77↓6.40
Experimental results indicate that using either isolated privacy keywords or the reference alone yields significantly lower performance than our approach. Removing the reference substantially impairs the model’s ability to generate valid questions, where utility’s Rouge-1 score is no more than 1%. Omitting private keywords disrupts the trade-off, resulting in slight privacy leakage that is disproportionate to the utility loss. Optimal performance is achieved only when both elements are combined. We attribute this to the fact that, although privacy keywords serve as suppression targets, LLMs can leverage these keywords to produce effective paraphrases, thereby enhancing utility.
7 Adversarial Attack
Based on the threat model in Section˜3, we design two adversarial experiments: Static and Adaptive attacks. In the static setting, the attacker is unaware of DP-GTR and directly attempts to recover the original question from cloud-received paraphrases. In the adaptive setting, the attacker fully understands DP-GTR, replicates its framework, and leverages privacy consensus keywords to infer the original question. In Table˜3, using GPT-3.5 and DocVQA as examples, we report the Rouge-1 (%) score between the original and DP-GTR-processed prompts under different releasing strategies, along with the corresponding $\Delta$ for each attack. See detailed attack settings in Appendix D.
The results of adversarial attacks demonstrate that DP-GTR offers strong robustness against both static and adaptive threats, with no observed increase in question Rouge-1 similarity, indicating no privacy leakage. These findings underscore effectiveness in safeguarding against adversarial risks.
@@ 8 Conclusion
This paper proposes DP-GTR, a novel three-stage local differential privacy (LDP) framework that leverages differentially private paraphrasing and the composition theorem through group text rewriting to enhance the privacy-utility trade-off. DP-GTR is the first approach, to our knowledge, to apply in-context learning for LDP prompt privatization and to connect global and local DP mechanisms via grouped paraphrased text. Furthermore, our framework is generalizable and compatible with any existing paraphrasing technique. Evaluations on open- and closed-answer QA datasets (DocVQA and Commonsense QA), simulating real-world LLM application scenarios, demonstrate that DP-GTR achieves a significantly superior privacy-utility trade-off compared to existing state-of-the-art methods. With the rapidly increasing adoption of LLMs, DP-GTR provides a practical, robust and readily deployable solution for mitigating the risk of user prompt privacy leakage.
@@ 9 Limitations
10 Acknowledgment
This work was supported in part by the CAHSI–Google Institutional Research Program (IRP) Grant and by the National Science Foundation (NSF) under grants CCF-2447834, HSI-2225229, CNS-2231519, and DUE-2225229.
Appendix A DP-guaranteed Paraphrasing Proof
Proof. Let $D$ and $D^{\prime}$ represent two arbitrary documents, and $\mathbf{u},\mathbf{u}^{\prime}\in\mathbb{R}^{|V|}$ denote the associated logit vectors. Set all the logits lie within the interval $b_{min},\,b_{max}]$ under sensitivity clipping. For a given token $v\in V$, let $i$ denote its position index and $u_{i}$ represent its corresponding logit from $\mathbf{u}$. Consequently, we obtain (Utpala et al., 2023),
$\begin{aligned} \frac{\text{Pr}[M(D)=v]}{\text{Pr}[M(D^{\prime})=v]}&=\frac{\frac{\exp(\frac{u_{i}}{T})}{\sum_{j=1}^{|V|}\exp(\frac{u_{j}}{T})}}{\frac{\exp(\frac{u_{i}^{\prime}}{T})}{\sum_{j=1}^{|V|}\exp(\frac{u_{j}^{\prime}}{T})}}\\ &=\frac{\exp(\frac{u_{i}}{T})}{\exp(\frac{u_{i}^{\prime}}{T})}\frac{\sum_{j=1}^{|V|}\exp(\frac{u_{j}^{\prime}}{T})}{\sum_{j=1}^{|V|}\exp(\frac{u_{j}}{T})}\\ &=\exp\left(\frac{u_{i}-u_{i}^{\prime}}{T}\right)\frac{\sum_{j=1}^{|V|}\exp(\frac{u_{j}^{\prime}}{T})}{\sum_{j=1}^{|V|}\exp(\frac{u_{j}}{T})}\\ &\leq\exp\left(\frac{b_{max}-b_{min}}{T}\right)\exp\left(\frac{b_{max}-b_{min}}{T}\right)\\ &\leq\exp\left(\frac{2(b_{max}-b_{min})}{T}\right).\end{aligned}$